OpenStudy (anonymous):

To solve the equation r divided by four minus two-ninths equals four-sixths for r, what two operations must be performed, and in which order? Add then multiply. Add then divide. Subtract then multiply. Subtract then divide.

OpenStudy (anonymous):

@RigelCanopus

OpenStudy (texaschic101):

(r/4) - (2/9) = 4/6 or is it ... r/(4-2/9) = 4/6

OpenStudy (anonymous):

the first one

OpenStudy (texaschic101):

(r/4) - 2/9 = 4/6 --- add 2/9 to both sides r/4 = 4/6 + 2/9 -- multiply both sides by 4 r = 4(4/6 + 2/9)

OpenStudy (texaschic101):

so your gonna add, then multiply